# Large-Scale Resilience Planning for Wildfire-Prone Electricity-System via Adaptive Robust Optimization

### 21 Mar 2026 v1

@@ III-A Planning and Operational Structure
Long-term grid planning. Utilities must first make long-term planning decisions regarding grid infrastructure and protection configuration, which determine the operational flexibility of the system during high-risk conditions. We consider two such planning decisions: feeder sectionalization and fast-trip protection configuration. First, feeder sectionalization divides large distribution circuits into smaller controllable segments [10, 11]. By installing sectionalizing devices, utilities can isolate portions of the network and selectively de-energize specific segments without shutting down the entire circuit, which improves operational flexibility. Second, utilities determine which sectionalized segments should be equipped with fast-trip protection [8, 9]. Fast-trip protection refers to protection settings that automatically disconnect a circuit when abnormal electrical conditions, such as faults or line contacts, are detected.
@@ III-B Tri-Level Optimization Model
$\min_{\mathbf{x},\mathbf{y}}\max_{\mathbf{u}\in\mathcal{U}}\min_{\mathbf{z}}\sum_{i=1}^{n}h_{i}(1-\alpha_{i}y_{i})(u_{i}-z_{i}),$ (1)
@@
The objective function in (1) captures the combined effects of operational and protection-based mitigation. First, PSPS actions are assumed to eliminate the electrical ignition source during the de-energization period [35, 36]. Thus, each PSPS action on segment $i$ removes one potential ignition event, so that applying $z_{i}$ actions reduces the number of ignition incidents from $u_{i}$ to $u_{i}-z_{i}$. Second, fast-trip protection reduces the likelihood that a fault escalates into a sustained ignition but does not fully eliminate it [37, 38]. When enabled ($y_{i}=1$), the expected number of ignition incidents is reduced by a segment-specific factor $\alpha_{i}$. Consequently, the residual ignition exposure contributing to system impact is modeled as $(1-\alpha_{i}y_{i})(u_{i}-z_{i})$.
@@ IV-B Uncertainty Set Construction
Conditional on $\mathcal{D}_{\mathrm{tr}}$, the bi-infinite score process $\{e_{k}\}_{k\in\mathcal{D}_{\mathrm{cal}}}$ is strictly stationary and $\alpha$-mixing. That is, its mixing coefficients
@@
Under Assumption 1, the uncertainty set $\mathcal{U}$ retains an approximate finite-sample coverage guarantee even when the calibration scores are temporally dependent.
@@
$\epsilon\coloneqq 12\Bigl(\frac{\Gamma}{2}\Bigr)^{1/3}\frac{(\log|\mathcal{D}_{\mathrm{cal}}|)^{2/3}}{|\mathcal{D}_{\mathrm{cal}}|^{1/3}}.$
@@
Proof.
Let $e$ denote the nonconformity score associated with the prediction over the future time period $T,T+\Delta)$. Under Assumption 1, Corollary 2 in [28] implies that
$\mathbb{P}\left(e\leq\hat{Q}(\alpha)~\middle|~\mathcal{D}_{\mathrm{tr}}\right)\geq 1-\alpha-\epsilon.$
By construction of the uncertainty set in (3) and the bounds in (5), the event $\{e\leq\hat{Q}(\alpha)\}$ is equivalent to $\{\mathbf{u}\in\mathcal{U}\}$. Therefore, this proves (6). ∎
@@ V-A Reformulation of the Lower-Level Recourse Problem
$\mathcal{P}(\mathbf{x},\mathbf{y},\mathbf{u})\coloneqq\min_{\mathbf{z},\mathbf{w}}$ $\sum_{i=1}^{n}h_{i}\big(u_{i}-z_{i}-\alpha_{i}u_{i}y_{i}+\alpha_{i}w_{i}\big)$ (8)
@@ V-C Mixed Integer Bi-level Subproblem
$\theta\leq\sum_{i=1}^{n}h_{i}\Big(u_{i}-z_{i}^{l}-\alpha_{i}u_{i}y_{i}^{m}+\alpha_{i}w_{i}^{l}\Big),$
@@ VI Synthetic Data Example
To reflect the temporal dependence observed in real weather data, we generate a scalar weather covariate $v_{k}$ shared across all segments from a stationary autoregressive process: $v_{k}=\rho v_{k-1}+\xi_{k}$, where $\xi_{k}\sim\mathcal{N}(0,\sigma^{2})$ and $|\rho|<1$ for stationarity. Heterogeneous weather sensitivity parameter $\beta_{i}$ are assigned to segments. Conditional on the scalar covariate $v_{ik}$ and noise term $\varepsilon_{ik}$, the ignition count on segment $i$ during interval $t_{k},t_{k+1}]$ is generated as $u_{ik}\sim\mathrm{Poisson}(\lambda_{ik})$, $\lambda_{ik}=\mu\exp(\beta_{i}v_{ik}+\varepsilon_{ik}),$ where $\mu$ controls the baseline ignition intensity. These simulated ignition counts are then used to train a linear Poisson model for segment- and group-level nonconformity scores.
@@ VII-A Dataset and Model Setup
Figure 3: Overview of the optimal sectionalization and fast-trip configuration in the California case study. The map highlights circuits where sectionalization and fast-trip protection are implemented across the study region, and they are concentrated in high-risk areas, broadly aligning with the CPUC-designated High Fire-Threat Districts (HFTD) [45].
To enrich the temporal dimension of the dataset, we construct a semi-synthetic data generator from the observed weather and ignition records. Following a common approach in meteorological research, we first reduce the dimension of the high-dimensional environmental features using principal component analysis (PCA) [46], and then fit a vector autoregressive (VAR) [47] model to the retained component scores to capture the dominant temporal dependence structure [48, 49]. We then simulate the fitted VAR forward to generate synthetic principal-component trajectories, map them back to the original weather-feature space, and convert the resulting environmental features into ignition outcomes using the fitted ignition model from the real data. Details are provided in Appendix -B.
@@
With the uncertainty set $\mathcal{U}$ and model parameters $W$, $B$, and $C$ specified, we implement Algorithm 1 in Python with Gurobi 11.0 [52]. Each outer iteration solves a MILP that yields candidate first-stage decisions, then runs an inner loop that alternates between an LP to identify the worst-case ignition distribution over $\mathcal{U}$ and a mixed-integer recourse MILP until convergence. We set $\varepsilon_{\mathrm{out}}=\varepsilon_{\mathrm{in}}=10^{-5}$.
@@ VII-B Results and Implications
Table I reports the case-study results comparing different baseline methods. We assess their out-of-sample performance from two perspectives: ($i$) We evaluate the empirical coverage of $\mathcal{U}$ over the planning period. ($ii$) We fix the first-stage decisions and recompute the optimal operational response to face the ignition in the held-out planning period, where $\|\mathbf{z}^{\text{true}}\|_{1}$ denotes the resulting number of PSPS actions. In addition to the baselines described in Section VI, we consider a Planning-Only benchmark that only solves the outer sectionalization and fast-trip planning problem in (1), and a Co-Optimized benchmark that removes the worst-case layer in (1) and solves a joint minimization problem for planning and recoure problems. For the proposed method, we implement it using randomly generated circuit groups, and we also report a variant based on the IOU’s official five-region service-territory partition as a fixed grouping structure with the same number of groups.
@@
Fig. 3 visualizes the optimal sectionalization and fast-trip configurations across the study region based on our method using real data (2020-2024). The recommended sectionalization actions are concentrated in the CPUC-designated High Fire-Threat Districts (HFTD) within the IOU’s service territory [45]. Fig. 4 further illustrates the effects of varying the planning parameters $C$ and $W$. As the sectionalization budget $C$ or the reliability limit $W$ increases, the model selects more sectionalized circuits, which provides greater operational flexibility to manage ignitions during the planning period through targeted PSPS or fast-trip actions.
@@ VII-C Managerial Insights
Table I also reveals several managerial insight. First, the results highlight the value of jointly optimizing long-term planning and short-term operations. Planning decisions that may appear less effective when evaluated in isolation can deliver substantially greater system benefits once their operational flexibility under severe wildfire conditions is taken into account. Second, the results suggest that sharper uncertainty quantification can meaningfully change the preferred mitigation strategy. In particular, tighter and valid uncertainty bounds tend to direct more investment toward network sectionalization, which improves operational flexibility, rather than relying primarily on broad preventive shutoffs or widespread fast-trip activation. For example, compared to the baselines, our method produces PSPS decisions that are much closer to those in the planning stage, whereas competing methods tend to overestimate the number of PSPS needed in the future period.

### 12 Apr 2026 v3

@@ III-A Planning and Operational Structure
Long-term grid planning. Utilities must first make long-term planning decisions regarding grid infrastructure and protection configuration, which determine the operational flexibility of the system during high-risk conditions. We consider two such planning decisions: feeder sectionalization and fast-trip protection configuration. First, feeder sectionalization divides large distribution feeders into smaller controllable segments [12, 2]. By installing sectionalizing devices, utilities can isolate portions of the network and selectively de-energize specific segments without shutting down the entire circuit, which improves operational flexibility. Second, utilities determine which sectionalized segments should be equipped with fast-trip protection [6, 9]. Fast-trip protection refers to protection settings that automatically disconnect a circuit when abnormal electrical conditions, such as faults or line contacts, are detected.
@@ III-B Tri-Level Optimization Model
$\min_{\mathbf{x},\mathbf{y}}\max_{\mathbf{u}\in\mathcal{U}}\min_{\mathbf{z}}\sum_{i=1}^{n}h_{i}(1-\beta_{i}y_{i})(u_{i}-z_{i}),$ (1)
@@
The objective function in (1) captures the combined effects of operational and protection-based mitigation. First, PSPS actions are assumed to eliminate the electrical ignition source during the de-energization period [27, 7]. Thus, each PSPS action on segment $i$ removes one potential ignition event, so that applying $z_{i}$ actions reduces the number of ignition incidents from $u_{i}$ to $u_{i}-z_{i}$. Second, fast-trip protection reduces the likelihood that a fault escalates into a sustained ignition but does not fully eliminate it [10, 31]. When enabled ($y_{i}=1$), the expected number of ignition incidents is reduced by a segment-specific factor $\beta_{i}$. Consequently, the residual ignition exposure contributing to system impact is modeled as $(1-\beta_{i}y_{i})(u_{i}-z_{i})$.
@@ IV-B Uncertainty Set Construction
Conditional on $\mathcal{D}_{\mathrm{tr}}$, the calibration nonconformity scores $\{e_{k}\}_{k\in\mathcal{D}_{\mathrm{cal}}}$ together with nonconformity score $e$ over the future time period $T,T+\Delta)$ form the observations of a strictly stationary $\alpha$-mixing process $\{e_{k}\}_{k\in\mathbb{Z}}$. That is, its mixing coefficients
@@
Under Assumption 1, the uncertainty set $\mathcal{U}$ retains a finite-sample coverage guarantee even when the calibration scores are temporally dependent.
@@
$\epsilon\coloneqq$ $2\Bigl(3+8\Gamma\Bigr)^{1/3}\frac{(3+\frac{\log|\mathcal{D}_{\mathrm{cal}}|}{2\log 2})^{2/3}}{|\mathcal{D}_{\mathrm{cal}}|^{1/3}}$ (7)
$=$ $\mathcal{O}\left(\frac{(\log|\mathcal{D}_{\mathrm{cal}}|)^{2/3}}{|\mathcal{D}_{\mathrm{cal}}|^{1/3}}\right).$
@@
The complete proof is deferred to Appendix -A. Theorem 1 shows that, under Assumption 1, the coverage gap $\epsilon$ decreases at the rate in (7). Hence, the resulting uncertainty set is asymptotically valid as the size of the calibration set grows.
@@ V-A Reformulation of the Lower-Level Recourse Problem
$\mathcal{P}(\mathbf{x},\mathbf{y},\mathbf{u})\coloneqq\min_{\mathbf{z},\mathbf{w}}$ $\sum_{i=1}^{n}h_{i}\big(u_{i}-z_{i}-\beta_{i}u_{i}y_{i}+\beta_{i}w_{i}\big)$ (9)
@@ V-C Mixed Integer Bi-level Subproblem
$\theta\leq\sum_{i=1}^{n}h_{i}\Big(u_{i}-z_{i}^{l}-\beta_{i}u_{i}y_{i}^{m}+\beta_{i}w_{i}^{l}\Big),$
@@ VI Synthetic Data Example
To reflect the temporal dependence observed in real weather data, we generate a scalar weather covariate $v_{k}$ shared across all segments from a stationary autoregressive process: $v_{k}=\rho v_{k-1}+\xi_{k}$, where $\xi_{k}\sim\mathcal{N}(0,\sigma^{2})$ and $|\rho|<1$ for stationarity. Heterogeneous weather sensitivity parameter $\kappa_{i}$ are assigned to segments. Conditional on the scalar covariate $v_{ik}$ and noise term $\varepsilon_{ik}$, the ignition count on segment $i$ during interval $t_{k},t_{k+1}]$ is generated as $u_{ik}\sim\mathrm{Poisson}(\lambda_{ik})$, $\lambda_{ik}=\mu\exp(\kappa_{i}v_{ik}+\varepsilon_{ik}),$ where $\mu$ controls the baseline ignition intensity. These simulated ignition counts are then used to train a linear Poisson model for segment- and group-level nonconformity scores. Details are provided in Appendix -B.
@@ VII-A Dataset and Model Setup
Figure 3: Overview of the optimal sectionalization and fast-trip configuration outcome in the California case study. The four panels compare the planning decisions produced by our method, a looser uncertainty set baseline (Bonferroni), Co-Optimized method, and Planning-Only method.
To enrich the temporal dimension of the dataset, we construct a semi-synthetic data generator from the observed weather and ignition records. Following a common approach in meteorological research, we first reduce the dimension of the high-dimensional environmental features using principal component analysis (PCA) [20], and then fit a vector autoregressive (VAR) [38] model to the retained component scores to capture the dominant temporal dependence structure [37, 49]. We then simulate the fitted VAR forward to generate synthetic principal-component trajectories, map them back to the original weather-feature space, and convert the resulting environmental features into ignition outcomes using the fitted ignition model from the real data. Details are provided in Appendix -C.
@@
With the uncertainty set $\mathcal{U}$ and parameters $W$, $B$, and $C$ fixed, we implement Algorithm 1 in Python with Gurobi 11.0 [17]. Each outer iteration solves a first-stage MILP, and each inner loop alternates between a worst-case LP over $\mathcal{U}$ and a recourse MILP until convergence. We set $\varepsilon_{\mathrm{out}}=\varepsilon_{\mathrm{in}}=10^{-5}$.
@@ VII-B Results and Implications
Table I reports the case-study results comparing different baseline methods. We assess their out-of-sample performance from two perspectives: ($i$) We evaluate the empirical coverage of $\mathcal{U}$ over the planning period. ($ii$) We fix the first-stage decisions given by each method, and recompute the optimal operational response to face the ignition in the held-out planning period by solving the recourse problem (9), where $\|\mathbf{z}^{\text{true}}\|_{1}$ denotes the resulting number of PSPS actions. We then report the corresponding cost obtained. In addition to the baselines described in Section VI, we consider a Planning-Only benchmark that only solves the outer sectionalization and fast-trip planning problem in (1), and a Co-Optimized benchmark that removes the worst-case layer in (1) and solves a joint minimization problem for planning and recourse problems. For the proposed method, we implement it using randomly generated circuit groups, and we also report a variant based on the IOU’s official five-region service-territory partition as a fixed grouping alternative with the same number of groups. Details of compared baselines are provided in Appendix -D.
@@
Fig. 3 provides a spatial view of the differences in Table I. Under our method, the selected sectionalization and fast-trip actions are concentrated in the CPUC-designated High Fire-Threat Districts (HFTD) within the IOU’s service territory [8], showing that the model prioritizes planning resources where operational flexibility is most valuable under wildfire risk. At the same time, our method recommends sectionalization without excessive reliance on fast-trip protection, instead addressing realized ignitions through adaptive operational actions, namely PSPS. By contrast, the baseline methods deploy fast-trip much more broadly over sectionalized circuits, reflecting either overly conservative risk estimates or less effective planning decisions. Accordingly, Table I shows that the Bonferroni and Max-Rank methods install roughly five times as many fast-trip configurations as our method.
Intuitively, this contrast follows from (1)–(2). Fast-trip protection $y_{i}$ lowers loss directly through $(1-\beta_{i}y_{i})$, so if $\mathcal{U}$ is overly conservative and inflates the worst-case value of $u_{i}$, the model tends to install fast-trip broadly ex ante. By contrast, sectionalization without fast-trip is valuable only when it enables adaptive PSPS and ignition risk materializes. To see this, for a sectionalized segment with $x_{i}=1$, marginal fast-trip reduces worst-case loss by approximately $h_{i}\beta_{i}u_{i}$, whereas PSPS reduces it by at most $h_{i}$. Thus, larger worst-case values of $u_{i}$ make fast-trip relatively more attractive. At the same time, fast-trip and PSPS share the same reliability budget in (2c), so a sharper uncertainty set makes it preferable to preserve that budget for targeted recourse rather than spend it on broad fast-trip deployment. This explains why our method, with a less conservative uncertainty set, favors more selective sectionalization over excessive fast-trip protection.
Fig. 4 further illustrates the effects of varying the planning parameters $C$ and $W$ with our method. As the sectionalization budget $C$ or the reliability limit $W$ increases, our model selects more sectionalized circuits, creating greater operational flexibility to manage ignitions during the planning period through targeted PSPS or fast-trip configuration.
@@ VII-C Managerial Insights
Table I and Fig 3 also reveal several managerial insight. First, the results highlight the value of jointly optimizing long-term planning and short-term operations. Planning decisions that may appear less effective when evaluated in isolation can deliver substantially greater system benefits once their operational flexibility under severe wildfire conditions is taken into account. Second, the results suggest that sharper uncertainty quantification can meaningfully change the preferred mitigation strategy. In particular, tighter and valid uncertainty bounds tend to direct more investment toward network sectionalization, which improves operational flexibility, rather than relying primarily on broad preventive shutoffs or widespread fast-trip activation. For example, compared to the baselines, our method produces PSPS decisions that are much closer to those in the planning stage, whereas competing methods tend to overestimate the number of PSPS needed in the future period.